The same heat transfer into identical masses of different substances produces different temperature changes. Calculate the final temperature when 1.00 kcal of heat transfers into 1.00 kg of the following, originally at 20.0
1
Expert's answer
2021-07-21T09:05:59-0400

a)


"T=20+\frac{1}{1(1)}=21\degree C"

b)


"T=20+\frac{1}{210(1)}=20.005\degree C"

c)


"T=20+\frac{1}{114(1)}=20.008\degree C"

d)


"T=20+\frac{1}{0.03(1)}=53\degree C"